OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[(\frac{ 1 }{ 3 }y^{3})^{2}\]

OpenStudy (anonymous):

\[(\frac{ 1 }{ 3 } )^2\] \[= \frac{ 1 }{ 9}\] \[(y^{3} )^{2} = y ^{3*2} = y ^{6}\] so the final simplified equation would be: \[\frac{ 1 }{ 9 } y ^{6}\]

OpenStudy (anonymous):

Thanks a lot! @ChihuahuaAdorer